OpenStudy (anonymous):

Need some guidance on this please... A tugboat goes 120 miles upstream in 20 hours. The return trip downstream takes 10 hours. Find the speed of the tugboat without a current and the speed of the current. The speed of the tugboat is ( ) mph, and the speed of the current is ( ) mph.

OpenStudy (dumbcow):

let v be speed of boat, and c be speed of current "upstream" is against current so subtract .... v-c "downstream" is with current so add ... v+c net speed is simply distance over time \[v-c = \frac{120}{20}\] \[v+c = \frac{120}{10}\] simplify and solve system using either substitution or elimination

OpenStudy (anonymous):

6 mph 12mph ??

OpenStudy (dumbcow):

yes, 120/20 = 6 and 120/10 = 12 but you still need to find v and c

OpenStudy (anonymous):

i am unsure how to do those steps, I am not sure what I am doing at all actually LOL

OpenStudy (dumbcow):

ok think of it this way.. what are 2 numbers which add to 12, but their difference is 6 ?

OpenStudy (anonymous):

6+6 is 12 4+8 is 12

OpenStudy (dumbcow):

"but their difference is 6"

OpenStudy (dumbcow):

6-6 = 0 8-4 = 4 no good, they have to subtract to 6

OpenStudy (anonymous):

12-6=6

OpenStudy (anonymous):

9-3=6

OpenStudy (anonymous):

3+9 =12

OpenStudy (anonymous):

and it's difference is 6

OpenStudy (anonymous):

Is that right 3+9=12 So 3mph and 9mph for V & C? Am I on the right track ?

OpenStudy (dumbcow):

you are correct the boat is going 9mph ...and then you either add or subtract the current of 3 mph

OpenStudy (anonymous):

A tugboat goes 120 miles upstream in 20 hours. The return trip downstream takes 10 hours. Find the speed of the tugboat without a current and the speed of the current. The speed of the tugboat is ( 9 ) mph, and the speed of the current is ( 12 ) mph. (i would add the 3 to find the current?) Can you check this.

OpenStudy (anonymous):

I think I confused myself somehow.

OpenStudy (anonymous):

Hope my answer is right